40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Jul 5 20:57:04 2019, Koji, Update, SUS, Another M7.1 EQ Screenshot_from_2019-07-05_21-03-06.png
    Reply  Fri Jul 5 21:53:10 2019, Koji, Update, SUS, Another M7.1 EQ 
       Reply  Fri Jul 5 22:21:13 2019, Koji, Update, SUS, Another M7.1 EQ Screenshot_from_2019-07-05_22-19-57.png
          Reply  Fri Jul 12 07:37:48 2019, gautam, Update, SUS, M4.9 EQ in Ridgecrest 56.pngITMXunstick.pngITMX_UL.pdf
             Reply  Tue Jul 16 15:00:03 2019, gautam, Update, SUS, Multiple small EQs 
Message ID: 14728     Entry time: Fri Jul 5 21:53:10 2019     In reply to: 14727     Reply to this: 14729
Author: Koji 
Type: Update 
Category: SUS 
Subject: Another M7.1 EQ 

- ITM unstuck now
- IMC briefly locked at TEM00

A series of aftershocks came. I could unstick ITMX by turning on the damping during one of the aftershocks.
Between the aftershocks, MC1~3 were aligned to the previous dof values. This allowed the IMC flashing. Once I got the lock of a low order TEM mode, it was easy to recover the alignment to have a weak TEM00.
Now at least temporarily the full alignment of the IMC was recovered.

ELOG V3.1.3-